# appendicesapc Appendix Table of Contents

### v1

@@ 3.2 Differentiable Responses Ranking
One of the most intuitive strategies to learn human preferences from the response list $\mathbf{y}$, is to sort the responses by predicted scores and use this ranking to fine-tune the language model, thereby learning the optimal preference ordering. However, traditional sorting methods like Selection Sort (Musser, 1997) and Quick Sort (Hoare, 1962) are inherently discrete and discontinuous, impeding differentiable optimization in LLM fine-tuning for preference learning (Petersen et al., 2021). In this section, we employ a differentiable sorting network [^2] Sorting networks are specialized computational architectures that sort sequences through comparisons and exchanges, not neural networks for sorting tasks. to rank responses based on scores $\hat{\mathbf{s}}_{\theta}$, enabling end-to-end fine-tuning of LLM on ranking preferences.
@@
Differentiable sorting networks offer superior parallel efficiency and excellent sorting performance while maintaining differentiability (Petersen et al., 2021). For a list of length $L$, the time complexity ranges from $\mathcal{O}(L^{2})$ to $\mathcal{O}(L\log^{2}{L})$, depending on the specific network variant (Akl, 1990) (such as Odd-Even or Bitonic networks) (Akl, 1990). These complexities are competitive with or surpass many differentiable sorting methods, including those proposed in (Song et al., 2024; Grover et al., 2019; Liu et al., 2024; Xia et al., 2008; Blondel et al., 2020; Swezey et al., 2021). A comprehensive time complexity comparison is provided in Table 5. Furthermore, sorting networks produce doubly stochastic permutation matrices, crucial for accurate NDCG computation by representing ranking probabilities faithfully. In contrast, existing differentiable sorting methods (Grover et al., 2019; Swezey et al., 2021) often produce unimodal permutation matrices, leading to overestimation in response gain calculations and severely distorting diffNDCG (see Eq. 13) measurements. The experimental results presented in Table 1 demonstrate that sorting networks significantly outperform existing differentiable sorting methods across various performance metrics.
@@
Given the predicted scores $\hat{\mathbf{s}}_{\theta}=\mathbf{M}_{\rm arp}(x,\mathbf{y};\pi_{\theta})=(\hat{s}_{1},\ldots,\hat{s}_{K})$ (as defined in Eq. 6), we employ a $K$-layer odd-even sorting network to sort these scores in descending order. This sorting network operates through a systematic alternation between odd and even indexed elements. Specifically, we select elements at odd and even indices in an alternating manner. For each selected element $\hat{s}_{j}$, we compare it with the subsequent element $\hat{s}_{j+1}$. If the elements are not in the desired descending order, we swap them to ensure that the larger score precedes the smaller one. This process is repeated across all $K$ layers until the entire sequence $\hat{\mathbf{s}}_{\theta}$ is sorted from highest to lowest.
Differentiable Swapping Operation. Generally, this swapping operation can be expressed as
$\hat{s}^{\prime}_{j}=\max(\hat{s}_{j},\hat{s}_{j+1}),~{}~{}~{}\hat{s}^{\prime}_{j+1}=\min(\hat{s}_{j},\hat{s}_{j+1}),~{}\forall j\in\{1,...,K-1\}.$
Since the operations of $\max$ and $\min$ are non-differentiable, we need to modify the swapping operation to ensure the ranking process is differentiable. Following (Petersen et al., 2021; 2022), we can refine the $\min$ and $\max$ as follows:
${\min}_{\rm soft}(\hat{s}_{j},\hat{s}_{j+1})$ $=\hat{s}_{j}\cdot h(\hat{s}_{j+1}-\hat{s}_{j})+\hat{s}_{j+1}\cdot(1-h(\hat{s}_{j+1}-\hat{s}_{j})),$ (7)
${\max}_{\rm soft}(\hat{s}_{j},\hat{s}_{j+1})$ $=\hat{s}_{j}\cdot(1-h(\hat{s}_{j+1}-\hat{s}_{j}))+\hat{s}_{j+1}\cdot h(\hat{s}_{j+1}-\hat{s}_{j}),$
where $h(\cdot)$ is a $s$-shaped function, which can be written as follows:
[…]
$\mathbf{P}_{k}={\rm diag}(\mathbf{p}_{1},\mathbf{p}_{3},\ldots)~{}~{}\text{or}~{}~{}{\rm diag}(1,\mathbf{p}_{2},\mathbf{p}_{4},\ldots),$ (9)
for the respective stages. By multiplying the permutation matrices from each layer, we construct the overall permutation matrix $\mathbf{P}_{\text{soft}}=\mathbf{P}_{1}\cdot\ldots\cdot\mathbf{P}_{K}$. The final sorted scores are given by $\hat{\mathbf{s}}_{\text{order}}=\mathbf{P}_{\text{soft}}^{\top}\hat{\mathbf{s}}_{\theta}$.
3.3 Differentiable Normalized Discounted Cumulative Gain Loss
To align human preferences, a direct strategy involves optimizing the cross-entropy loss between the ground truth permutation matrix $\mathbf{P_{\rm ground}}$ (obtained by ground truth scores $\mathbf{s}$) and the predicted soft permutation matrix $\mathbf{P}_{\rm soft}$: let $\ell_{\rm ce}$ be the cross-entropy loss and $[\mathbf{P}]_{j}$ be the $j$-th column of matrix $\mathbf{P}$,
However, experiments in Table 4 reveals this strategy’s suboptimality. One possible explanation is that it fails to distinguish error severity across ranking positions, incorrectly equating misplacements of top-ranked responses with lower-ranked items, despite higher-ranked responses typically being far more crucial. To address these challenges, we propose optimizing Normalized Discounted Cumulative Gain (NDCG) (Järvelin and Kekäläinen, 2002b), an effective LTR metric for measuring ranking quality. NDCG assesses the significance of responses in conjunction with their ranking positions. It assigns greater importance to responses at the top of the ranking compared to those positioned lower, and imposes a more severe penalty for inaccurately placing a top-ranked response in a lower position.
Furthermore, while NDCG was originally designed to reflect users’ tendency to focus on top-ranked results (Järvelin and Kekäläinen, 2002a), this characteristic aligns with human preference rankings, which prioritize more preferred responses over less preferred ones (Pool et al., 2016). This similarity makes NDCG an effective proxy for evaluating and learning human preferences. Experiments in Figure 4 demonstrate that has a stronger correlation with human preference win rates compared to the optimizing targets of existing methods, highlighting its effectiveness. Additionally, NDCG prioritizes top-ranked responses and penalizes their misplacement, and capturing graded importance between responses without introducing any additional computation burden. Our time complexity analysis in Table 4 shows our method has comparable computational complexity to other methods.
$\mathrm{NDCG}(\hat{\mathbf{s}}_{\theta},\mathbf{s})$ $=\frac{1}{\rm iDCG}\sum_{j=1}^{K}\frac{2^{s_{j}}-1}{\log_{2}(1+q(y_{j}))},~{}\text{where}~{}{\rm iDCG}=\sum_{j=1}^{K}\frac{{2^{s_{j}}-1}}{{\log_{2}(1+q^{*}(y_{j}))}},$ (11)
@@ 3.3 Differentiable Normalized Discounted Cumulative Gain Loss
here $q(y_{j})$ is the ranking position of $y_{j}$ with respect to $\hat{\mathbf{s}}_{\theta}$ and $q^{*}(y_{j})$ is the ranking position of $y_{j}$ with respect to $\mathbf{s}$. NDCG assigns gains of $12^{s_{j}}-$ 1 based on the relevance score $s_{j}$ of each response $y_{j}$. It also applies discount factors ${\log_{2}(1+q(y_{j}))}$, where $q(y_{j})$ is the ranking position of $y_{j}$. This discounting mechanism assigns higher weights to elements at the top of the ranking. Consequently, NDCG effectively accounts for the varying importance of responses at different ranking positions.
However, the computation method for NDCG involves a sorting process that renders the metric non-differentiable with respect to the ranking position $q(y_{j})$.
$\mathrm{NDCG}(\hat{\mathbf{s}}_{\theta},\mathbf{s})$ $=\frac{1}{\rm iDCG}\sum_{j=1}^{K}\frac{2^{\psi(d,\mathbf{s},\hat{\mathbf{s}}_{\theta})}-1}{\log_{2}(1+d)}=\frac{1}{\rm iDCG}\sum_{d=1}^{K}\frac{2^{\psi(d,\mathbf{s},\hat{\mathbf{s}}_{\theta})}-1}{\log_{2}(1+d)}.$ (12)
[…]
Finally, we consider the following optimization problem:
$\min_{\theta}\mathcal{L}_{\rm diffNDCG}=\frac{1}{|D|}\sum_{(x,\mathbf{y},\mathbf{s})\in D}\ell_{\rm diffNDCG}(\mathbf{M}_{\rm ad}(x,\mathbf{y},\pi_{\theta}),\mathbf{s}),$ (14)
where $\ell_{\rm diffNDCG}(\cdot,\cdot)=-\mathrm{diffNDCG}(\cdot,\cdot).$ The pseudo code of DRPO is presented in Appendix C.
4 Experiments
4.1 Experimental Settings
Datasets. Anthropic’s Helpful and Harmless (HH) (Bai et al., 2022) contains 161k/8.5k training/test samples. Each sample consists of a prompt and a pair of responses (chosen and reject), where “chosen" represents the preferred response and “reject" represents the less preferred response. We also generate additional responses for each prompt and rate each response using a reward model DeBERTa  [^3] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2 , resulting in ranking preference data of a list size $K=8$. For more details, please refer to Appendix D. UltraFeedback (Cui et al., 2023) contains 64k prompts. Each prompt corresponds to four responses, and every response has a score annotated by AI (e.g., GPT-4 and Gemini). VLFeedback (Li et al., 2023a) consists of 80k multi-modal samples from various sources. Each sample contains four responses from different models and is annotated by GPT-4V (OpenAI, 2023).
Models. Our experiments are mainly based on Qwen1.5 model (Bai et al., 2023a) with a range of parameters from 0.5B to 1.8B and Mistral model (Jiang et al., 2023) with 7B parameter size. In addition, we also train Qwen-VL-Chat (Bai et al., 2023b), a large-scale vision-language model, to evaluate the performance of our method on multi-modal preference alignment task (Sun et al., 2023). A comprehensive experimental setup is provided in Appendix E.
Baseline Methods. To validate the effectiveness of our method, we conduct comparison experiments with representative baselines. In our experiments, we mainly compare our method with SFT, DPO (Rafailov et al., 2023), PRO (Song et al., 2024), LiPO (Liu et al., 2024), DPOBT and DPOPL (Rafailov et al., 2023). DPOBT adapts DPO to ranking preferences by decomposing ranked lists into pairwise comparisons, as proposed by Liu et al. (2024). Furthermore, DPOPL is proposed based on ranking preference data by (Rafailov et al., 2023), which leverages the Plackett-Luce preference model (Plackett, 1975), a generalization of the Bradley-Terry model (Bradley and Terry, 1952) that accommodates full rankings rather than just pairwise comparisons.
Furthermore, based on our previously introduced Policy Reference Ratio Score (PRR) and Adaptive Rank Policy Score (ARP) in Section 3.1, we implemented several differentiable sorting algorithms such as Fast Soft Sort (Blondel et al., 2020), Neural Sort (Grover et al., 2019) and learning-to-rank methods, including ListNet (Xia et al., 2008), PiRank (Swezey et al., 2021) for preference alignment.
Table 1: Comparisons between our method and competitors on Anthropic’s Helpful and Harmless Dataset. We report GPT-4 Win Rate (vs Chosen) and Reward Model Win Rate (vs Chosen and SFT).
Base Model Qwen1.5-0.5B Qwen1.5-1.8B
[…]
Diff Sorting Network 28.97%($\pm$ 1.46) 44.53%($\pm$ 1.65) 66.41%($\pm$ 3.78) 49.39%($\pm$ 2.55) 64.06%($\pm$ 1.75) 78.56%($\pm$ 2.31)
  • GPT-4 based Win Rate: we use GPT-4 to compare which of generated responses is more preferred, and evaluate the method’s performance by calculating the win rate of response.
  • Open Benchmarks: we evaluate different method on the UltraFeedback dataset using the AlpacaEval2.0 (Li et al., 2023b) and MT-Bench (Zheng et al., 2023) benchmarks. For the VLFeedback dataset, we use the MME (Fu et al., 2023), MM-Bench (Liu et al., 2023b), and MM-Vet (Yu et al., 2023) benchmarks to evaluate the performance.
4.2 Experimental Results on Anthropic’s Helpful and Harmless Dataset
Main Results. Experiments on HH dataset are conducted in Table 1 showing the effectiveness of our method. (1) Our method DRPO outperforms baselines SFT, DPOBT, DPOPL, PRO, LiPO and other LTR methods across different model scales. The GPT-4 Win Rate has an improvement of $5.22\%{\sim}6.13\%$, and Reward Model Win Rate has an improvement of $4.69\%{\sim}8.98\%$ and $0.78\%{\sim}2.73\%$. (2) Even without additional modifications, our Sorting Network outperforms most conventional sorting methods, such as Neural Sort. (3) Directly extending the existing alignment method, such as DPO, to ranking list scenarios can help improve performance. (4) LiPO employs a metric weighting scheme to account for the relative importance of responses at different positions, resulting in superior performance compared to DPOBT. In contrast, DRPO leverages diffNDCG to precisely quantify response contributions at each ranking position, significantly enhancing its performance over LiPO. Additionally, Appendix G.1 compares reward distributions of model-generated responses, revealing significant improvements in response rewards when using our method.
Table 2: A comprehensive analysis of integrating Adaptive Rank Policy Score into existing methods to enhance model performance.
[…]
Qwen1.5-0.5B
DPOBT(K=2) 28.68% 39.45% 62.89%
DPOBT(K=2) + ARP 31.80%(+3.12) 43.55%(+4.10) 68.16%(+5.27)
PiRank 26.08% 38.87% 62.50%
PiRank + ARP 42.35%(+16.27) 57.03%(+18.16) 77.14%(+14.64)
Fast Soft Sort 37.58% 51.95% 74.99%
Fast Soft Sort + ARP 40.69%(+3.11) 54.49%(+2.54) 77.92%(+2.93)
DRPO 42.80% 79.88% 79.88%
Qwen1.5-1.8B
DPOBT(K=2) 49.26% 56.25% 71.49%
DPOBT(K=2) + ARP 50.43%(+1.17) 59.57%(+3.32) 74.41%(+2.92)
PiRank 50.71% 56.64% 69.14%
PiRank + ARP 66.56%(+15.85) 75.69%(+18.95) 87.10%(+17.96)
Fast Soft Sort 61.90% 71.87% 85.93%
Fast Soft Sort + ARP 66.14%(+4.24) 77.92%(+6.05) 86.91%(+0.98)
DRPO 69.08% 82.61% 89.06%
Table 3: Effectiveness analysis of each component in our method.
[…]
DRPO-w/o ARP & diffNDCG 28.97%($\pm$1.46) 44.53%($\pm$1.65)
DRPO-w/o diffNDCG 33.06%($\pm$5.33) 47.66%($\pm$3.49)
DRPO-w/o ARP 38.30%($\pm$3.35) 53.91%($\pm$1.46)
DRPO 42.80%($\pm$5.01) 58.40%($\pm$2.94)
Table 4: Impact of different discount factors on diffNDCG performance.
[…]
$1/{\sqrt{r}}$ 40.53%($\pm$4.36) 79.29%($\pm$2.89)
$1/r$ 43.37%($\pm$5.14) 78.32%($\pm$2.94)
$1/{r^{2}}$ 40.37%($\pm$4.17) 79.30%($\pm$3.38)
$1/{\mathrm{log}(1+r)}$ 42.80%($\pm$5.01) 79.88%($\pm$3.72)
Ablation Studies on Adaptive Rank Policy Score. To assess the efficacy of our proposed Adaptive Rank Policy Score, we conducted a comparative analysis by integrating it into multiple baseline models. The results of our analysis are presented in Table 4. Our experimental results demonstrate that the Adaptive Rank Policy Score significantly enhances baseline model performance across various sizes, showcasing the advantages of applying ranking position-based adaptive margins over the PRR score.
Figure 3: Comparisons between DRPO and other methods across different list sizes on HH dataset.
Furthermore, we conduct an ablation study on our method by progressively removing Adaptive Rank Policy Score(ARP) and diffNDCG, denoted as ’w/o ARP’ and ’w/o diffNDCG’ respectively. As shown in Table 4, we observe that directly using the cross-entropy loss between the ground truth permutation matrix $\mathbf{P}_{\rm ground}$ and the predicted permutation matrix $\mathbf{P}_{\rm soft}$, without ARP and diffNDCG, leads to performance degradation. This fails to account for the varying importance of responses at different ranking positions. Additionally, when the ARP is replaced by PRR in Eq (2), performance decreases by 4.09-4.5% and 3.13-4.49% for GPT-4 Win Rate and RM Win Rate, respectively.
Table 5: Time and memory complexity comparisons with other methods.
Method Time Complexity Run Time Memory
PRO $\mathcal{O}(L^{2})$ 0.2806s 24.57GB
LiPO $\mathcal{O}(L^{2})$ 0.3269s 24.08GB
Neural Sort $\mathcal{O}(L^{2})$ 0.2535s 24.01GB
Pirank $\mathcal{O}(L^{2})$ 0.2535s 25.80GB
Fast Soft Sort $\mathcal{O}(L\log(L))$ 0.2618s 23.21GB
ListNet $\mathcal{O}(L^{2})$ 0.2641s 21.10GB
DRPO (odd-even) $\mathcal{O}(L^{2})$ 0.2641s 23.39GB
DRPO (bitonic) $\mathcal{O}(L^{2})$ 0.2560s 23.43GB
Ablation Studies on DiffNDCG Discounts Factors. In addition, we conducted experiments with different discount factors in diffNDCG using Qwen1.5-0.5B, details in Appendix G.2. The results are in Table 4. We discovered that various discount factors are indeed effective. Among them, the inverse log discount emerged as a well-balanced choice, offering an effective compromise between emphasizing top-ranked responses and penalizing their misplacements.
Ablation Studies on List Size. We also analyze the performance for different list sizes $K$ of ranking preference data. We compared various list-wise methods and included DPO as a comparison baseline on the HH dataset, using Qwen1.5-0.5B as the base model. As Figure 3 shows, our method almost always outperforms others across different list sizes $K$. Additionally, all list-wise methods, consistently outperform DPO, even at $K=2$, highlighting the advantage of list-wise over pair-wise approaches.
Figure 4: Relationships among Win Rate, NDCG, and Accuracy during training for two models: Qwen 1.5-0.5B (blue) and Qwen 1.5-1.8B (orange) when using DRPO. (a) NDCG versus RM Win Rate, (b) Ranking Accuracy versus RM Win Rate, and (c) NDCG versus Ranking Accuracy.
Time Complexity Analysis And Computational Efficiency. Given a list of length $L$, our differentiable sorting network achieves a time complexity of $\mathcal{O}(1)$ for each swap operation. When using an odd-even sorting network, the overall time complexity for differentiable sorting is $\mathcal{O}(L^{2})$. Additionally, computing the diffNDCG also has a complexity of $\mathcal{O}(L^{2})$. Consequently, the final time complexity of our algorithm is $\mathcal{O}(L^{2})$, which matches most differentiable sorting methods, such as Neural Sort, PiRank, LiPO. In the Table 5, we present a comparison of the time complexities and actual running times of different sorting algorithms. Here, the running time refers to the average time taken to compute the loss using the qwen1.5-0.5b model on 256 samples. With the maximum response list length set to a small value $L=8$, the algorithmic complexity across methods is similar.
Relationships Between Win Rate, NDCG, and Ranking Accuracy. To analyze the correlation between NDCG and evaluation metrics such as win rates, we extracted checkpoints during the training and quantified multiple metrics, including NDCG, RM Win Rate, and Ranking Accuracy. The experimental results are illustrated in Figure 4. We also directly measured the agreement between the metrics in Appendix G.3. Experimental results demonstrate a strong correlation between NDCG and evaluation metrics such as win rate and ranking accuracy. This high level of consistency suggests that optimizing NDCG can effectively improve both win rate and ranking accuracy in evaluation.
4.3 Experimental Results on UltraFeedback and VLFeedback Dataset
Table 6: Comparative analysis of aligning Mistral-7B-Base on the UltraFeedback dataset.
SFT 6.3 8.4% 6.2%
DPO 7.3 15.1% 12.5%
ORPO 7.3 14.7% 12.2%
R-DPO 7.4 17.4% 12.8%
LiPO 7.4 25.3% 18.9%
DRPO 7.3 26.5% 19.6%
Table 7: Comparisons between our method and baselines on multi-modal VLFeedback dataset.
Model MMEP $\uparrow$ MM-Bench $\uparrow$ MM-Vet $\uparrow$
DPO 1496.7 52.83% 45.2
DPOBT 1548.1 52.55% 46.8
LiPO 1559.3 54.55% 47.2
DRPO 1581.1 56.19% 48.6
Main Results on UltraFeedback Dataset. To validate scalability and performance, we train a Mistral model on the UltraFeedback dataset and evaluate it using open benchmarks. As shown in Table 7, our method can scale up to larger model and outperform other methods. In MT-Bench, our performance is comparable to existing methods. we achieve improvements of 1.2% in AlpacaEval2.0 length-controlled Win Rate (LC) and 0.7% in AlpacaEval2.0 raw Win Rate (WR).
Main Results on Multi-Modal VLFeedback Dataset. We aslo apply DRPO to fine-tune vision-language models. We use Qwen-VL-Chat as our base model and evaluate each method with multi-modal benchmarks. As Table 7 shows, our method outperform other methods. In MME benchmark, our method outperforms others by $21.8$ in perception tasks. Meanwhile, we achieve an improvement of $1.4$ in the MM-Vet benchmark and an improvement of $1.64\%$ in MM-Bench.
5 Conclusion
Aligning LLMs with human preferences is crucial for enhancing human-machine interactions and ensuring the safety of LLMs. However, most existing methods focus only on alignment with pairwise preference data, categorizing responses into preferred and less preferred pairs. Our research provides timely insights on how to align LLMs with multiple responses. We propose a novel method DRPO, which treats human preference alignment as a listwise ranking problem and aligns LLMs using ranking preference data. Specifically, we introduce a new ranking score computation strategy, termed Adaptive Rank Policy Score. Furthermore, we utilize the NDCG metric to develop our diffNDCG loss function. Extensive experiments demonstrate the superiority of our proposed method. It not only improves alignment performance but also opens avenues for future explorations in the field.
Appendix A Related Works
Human Preference Alignment. Although LLMs have demonstrated impressive capabilities across a broad range of tasks, they sometimes produce harmful and offensive content. This tendency may lead to the AI system’s insecurity and loss of control. To address this issue, Reinforcement Learning with Human Feedback (RLHF) was proposed by Bai et al. [2022], Christiano et al. [2017] to align LLMs with human preferences. This method involves first training a reward model using a preference model such as the Bradley-Terry model [David, 1963], and then fine-tuning LLMs with the trained reward model to maximize the given reward using reinforcement learning algorithms like PPO [Schulman et al., 2017]. However, RLHF faces challenges such as instability and high computational cost, which render the fine-tuning of LLMs using RLHF particularly challenging.
To alleviate the reliance on reinforcement learning, various methodologies have been proposed [Yuan et al., 2023, Hong et al., 2024, Ethayarajh et al., 2024]. For example, Direct Preference Optimization (DPO) [Rafailov et al., 2023] enables the extraction of the optimal policy in a closed form and derives a pairwise logistic loss directly from pairwise preferences. DPO eliminates the need for explicit reward modeling or reinforcement learning, thereby reducing the training costs associated with RLHF. Liu et al. [2023a] propose RSO, which utilizes the rejection sampling method to source preference data from the estimated target optimal policy, leading to a more accurate estimation of the optimal policy. They also propose a unified framework that enhances the loss functions used in both SLiC [Zhao et al., 2023] and DPO from a preference modeling standpoint.
However, most recent studies [Cheng et al., 2023, Azar et al., 2024, Yuan et al., 2023] only focus on pairwise preference cases, while only a few works focus on list preferences. RRHF, as proposed by Yuan et al. [2023], employs a pairwise hinge loss to align LLMs with collected ranking preference data. However, it treats each pair within the list as independent comparisons, rather than considering the overall ranking preference structure. PRO [Song et al., 2024] learns from listwise preference data using the list MLE loss, which based on the Plackett-Luce model [Plackett, 1975] rather than Bradley-Terry model. LiPO [Liu et al., 2024] proposes the first list preference optimization framework that conceptualizes human preference alignment as a learning-to-rank problem and unifies existing pairwise preference optimization methods into their framework. Additionally, they utilize a weighted pairwise logistic loss (lambda loss) [Wang et al., 2018] to align human preferences.
Utilizing ranking preference data to address human preference alignment is still a challenging task that has been scarcely explored to date. In this work, we propose diffNDCG, a novel method based on entire ranking preference data, which can utilize list data more effectively than existing methods.
Learning-to-Rank Task. Learning-to-Rank (LTR) task is a well-studied field with extensive literature, primarily due to its practical applications in web search and recommendation systems [Liu et al., 2009, Dai et al., 2011, Macdonald et al., 2013, Cao et al., 2007, Yu et al., 2013, 2019, Abdollahpouri et al., 2017, Zhang et al., 2024]. Traditional research in Learning to Rank (LTR) has concentrated on developing robust ranking objectives, including pointwise, pairwise, and listwise approaches. RankSVM [Cao et al., 2006] and RankNet [Burges et al., 2005] utilize pairwise hinge loss and pairwise logistic loss, respectively, to optimize ranking performance. ListMLE and Softmax losses are two representative listwise losses introduced by Cao et al. [2007]. LambdaRank [Burges et al., 2006] employs a pairwise logistic loss with lambda weights and achieves strong empirical performance improvements compared to RankNet.
Furthermore, numerous methods have been proposed to directly optimize the non-smooth NDCG metric. For instance, SoftRank [Taylor et al., 2008] employs rank distributions to smooth NDCG. ApproxNDCG [Qin et al., 2010a] uses a generalized sigmoid function to approximate the indicator function for rank computation and the top-K selector for the top-K variant. PiRank [Swezey et al., 2021] and NeuralNDCG [Pobrotyn and Bialobrzeski, 2021] approximate the non-continuous sorting operator using NeuralSort [Grover et al., 2019] to smooth NDCG. Inspired by these studies, we utilize differentiable sorting networks to smooth NDCG and introduce the diffNDCG loss function.
Appendix B Limitations
In conducting our experiments, we adhere to benchmarks that are widely recognized and routinely utilized in the literature on human preference alignment. It is important to note, however, that despite our best efforts, there remain certain unavoidable limitations that merit consideration. In the process of constructing a ranking preference dataset, we employ a reward model to serve as a proxy for human evaluations. While the reward model we employ achieves a high level of accuracy, it is important to acknowledge that there might still be subtle discrepancies when compared to human evaluations. In the future, we are also interested in exploring more high-quality reward models as proxies for human evaluations.
Appendix C DRPO Algorithm
@@
2:  Initialize: Policy $\pi_{\theta}$, reference policy $\pi_{\rm ref}$

### v2

@@ IV-B Differentiable Responses Ranking
One of the most intuitive strategies to learn human preferences from the response list $\mathbf{y}$, is to sort the responses by predicted scores and use this ranking to fine-tune the language model, thereby learning the optimal preference ordering. However, traditional sorting methods (e.g. Selection Sort [54], Quick Sort [55]) operate in a discontinuous manner, impeding differentiable optimization in LLM fine-tuning for preference learning [24]. In this section, we employ a differentiable sorting network [^1] Sorting networks are specialized computational architectures that sort sequences through comparisons and exchanges, not neural networks for sorting tasks. to rank responses based on scores $\hat{\mathbf{s}}_{\theta}$, enabling end-to-end fine-tuning of LLM on ranking preferences.
@@
Differentiable sorting networks offer superior parallel efficiency and excellent sorting performance while maintaining differentiability [24]. For a list of length $L$, the time complexity ranges from $\mathcal{O}(L^{2})$ to $\mathcal{O}(L\log^{2}{L})$, depending on the specific network variant [56] (such as Odd-Even or Bitonic networks) [56]. These complexities are competitive with or surpass many differentiable sorting methods, including those proposed in [21, 47, 15, 57, 58, 46]. A comprehensive time complexity comparison is provided in Table VII. Furthermore, sorting networks produce doubly stochastic permutation matrices, crucial for accurate NDCG computation by representing ranking probabilities faithfully. In contrast, existing differentiable sorting methods [47, 46] often produce unimodal permutation matrices, leading to overestimation in response gain calculations. The experimental results in Table I demonstrate that sorting networks significantly outperform existing differentiable sorting methods in various metrics.
Odd-even Sorting Network. In this work, we adopt the odd-even sorting network [59] for response ranking, which shows promising performance in our experiments while maintaining implementation simplicity and efficiency. Given the predicted scores $\hat{\mathbf{s}}_{\theta}=\mathbf{M}_{\rm arp}(x,\mathbf{y};\pi_{\theta})=(\hat{s}_{1},\ldots,\hat{s}_{K})$ (as defined in Eq. 6), we employ a $K$-layer odd-even sorting network to sort these scores in descending order. This sorting network alternates between odd and even indices. In each layer, selected elements $\hat{s}_{j}$ are compared with their subsequent elements $\hat{s}_{j+1}$ and swapped if not in descending order. This process continues through $K$ layers until sequence $\hat{\mathbf{s}}_{\theta}$ is sorted. Appendix presents a comprehensive procedure of this sorting network.
Differentiable Swapping Operation. Generally, For adjacent pairs $(j,j+1)$ where $j\in{1,...,K-1}$, the swapping operation is defined as:
$\hat{s}^{\prime}_{j}=\max(\hat{s}_{j},\hat{s}_{j+1}),~~~\hat{s}^{\prime}_{j+1}=\min(\hat{s}_{j},\hat{s}_{j+1}),$
Since the operations of $\max$ and $\min$ are non-differentiable, we need to modify the swapping operation to ensure the ranking process is differentiable. Following [24, 60], we can refine the $\min$ and $\max$ as follows:
${\min}_{\rm soft}(\hat{s}_{j},\hat{s}_{j+1})$ $=\hat{s}_{j}\cdot\alpha+\hat{s}_{j+1}\cdot(1-\alpha),$ (7)
${\max}_{\rm soft}(\hat{s}_{j},\hat{s}_{j+1})$ $=\hat{s}_{j}\cdot(1-\alpha)+\hat{s}_{j+1}\cdot\alpha,$
$\alpha$ $=h(\hat{s}_{j+1}-\hat{s}_{j}),$
where $h(\cdot)$ is a $s$-shaped function, which can be defined as:
here $\alpha$ represents the steepness that controls the relaxation strength. Then, we can reformulate the differentiable swapping operations described in Eq. 7 for the $k\text{-th}$ layer using a $K\times K$ permutation matrix, denoted as $\mathbf{P}_{k}$. Specifically, a swapping operation at index $j$ during either the odd or even stage can be represented as $\hat{s}^{\prime}_{j},\hat{s}^{\prime}_{j+1}]=[\hat{s}_{j},\hat{s}_{j+1}]\cdot\mathbf{p}_{j}$, where $\mathbf{p}_{j}$ is a $2\times 2$ matrix defined as:
To integrate all swapping operations in the $k\text{-th}$ layer, we aggregate the matrices corresponding to all odd or even indices, yielding the permutation matrix of the $k$-th layer.
$\mathbf{P}_{k}={\rm diag}(\mathbf{p}_{1},\mathbf{p}_{3},\ldots)~~\text{or}~~{\rm diag}(1,\mathbf{p}_{2},\mathbf{p}_{4},\ldots).$ (9)
Multiplying the permutation matrices of all layers yields the overall permutation matrix $\mathbf{P}_{\text{soft}}=\mathbf{P}_{1}\cdot\ldots\cdot\mathbf{P}_{K}$. The final ordered scores are given by $\hat{\mathbf{s}}_{\text{order}}=\mathbf{P}_{\text{soft}}^{\top}\cdot\hat{\mathbf{s}}_{\theta}$.
IV-C Differentiable Normalized Discounted Cumulative Gain Loss
To align human preferences with differentiable sorting, a direct strategy involves optimizing the cross-entropy loss between the ground truth permutation matrix $\mathbf{P_{\rm ground}}$ (obtained by ground truth scores $\mathbf{s}$) and the predicted soft permutation matrix $\mathbf{P}_{\rm soft}$: let $\ell_{\rm ce}$ be the cross-entropy loss and $\mathbf{P}]_{j}$ be the $j$-th column of matrix $\mathbf{P}$,
However, experiments in Table V reveals this strategy’s suboptimality. One possible explanation is that it fails to distinguish error severity across ranking positions, incorrectly equating misplacements of top-ranked responses with lower-ranked items, despite higher-ranked responses typically being far more crucial. To address these challenges, we propose optimizing Normalized Discounted Cumulative Gain (NDCG) [61], an effective LTR metric for measuring ranking quality. NDCG assesses the significance of responses in conjunction with their ranking positions. It assigns greater importance to responses at the top of the ranking compared to those positioned lower, and imposes a more severe penalty for inaccurately placing a top-ranked response in a lower position.
Furthermore, while NDCG was originally designed to reflect users’ tendency to focus on top-ranked results [61], this characteristic aligns with human preference rankings, which prioritize more preferred responses over less preferred ones [62]. This similarity makes NDCG an effective proxy for evaluating and learning human preferences. Experiments in Figure LABEL:fig:ndcg-winrate demonstrate that has a stronger correlation with human preference win rates compared to the optimizing targets of existing methods, highlighting its effectiveness. Additionally, NDCG prioritizes top-ranked responses and penalizes their misplacement, and capturing graded importance between responses without introducing any additional computation burden. Table VII shows our method achieves comparable computational complexity to existing methods.
$\mathrm{NDCG}(\hat{\mathbf{s}}_{\theta},\mathbf{s})$ $=\frac{1}{\rm iDCG}\sum_{j=1}^{K}\frac{2^{s_{j}}-1}{\log_{2}(1+q(y_{j}))},$ (11)
@@ IV-C Differentiable Normalized Discounted Cumulative Gain Loss
Here $q(y_{j})$ and $q^{*}(y_{j})$ denote the ranking positions of $y_{j}$ with respect to $\hat{\mathbf{s}}_{\theta}$ and $\mathbf{s}$, respectively. NDCG assigns gains of $2^{s_{j}}-1$ based on the score $s_{j}$ of each response $y_{j}$. It also applies discount factors $\log_{2}(1+q(y_{j}))$} to each response according to its rank $q(y_{j})$. This discount assigns higher weights to elements at the top of the ranking. Consequently, NDCG effectively accounts for the varying importance of responses at different ranking positions. However, the required sorting operation in NDCG makes it non-differentiable with respect to ranking positions $q(y_{j})$.
@@
2:Initialize: Policy $\pi_{\theta}$, reference policy $\pi_{\rm ref}$ .